# A Network-Based Measure of Cosponsorship Influence on Bill Passing in the United States House of Representatives

### v1

@@ Appendix A Appendix
In this appendix, we explore the influence of bill cosponsorship networks on bill-passage success using the standard centrality notions of closeness centrality (see Figure 5) and strength centrality (i.e., weighted degree centrality) (see Figure 6).
To compute closeness centrality and strength centrality [18], we use the largest connected component of a month-$t$ Congressional cosponsorship network and its corresponding adjacency matrix $D[t]_{\text{LCC}}$ (see Section 5.3). The closeness centrality of representative $i$ at month $t$ is
@@
where $N_{{\text{LCC}}}$ is the number of nodes in $D[t]_{\text{LCC}}$ and the entry $d_{ji}$ of $D[t]_{\text{LCC}}$ is the shortest-path distance from node $j$ to node $i$. The strength centrality of representative $i$ at month $t$ is
@@
Figure 5: Closeness-centrality time series and the corresponding distributions of relative differences between the closeness centralities of bills that pass and fail in the House of Representatives. In (a), we show the bill-level closeness centrality that we calculate using the mean of the closeness centralities of a bill’s cosponsors. In (b), we show the bill-level closeness centrality that we calculate using the maximum of the closeness centralities of a bill’s cosponsors. In the left panels, we show bill-level closeness centrality for 10 years, which we aggregate into nonoverlapping 4-month periods. Each point is the mean closeness centrality of bills in the associated 4-month period starting on the date on the horizontal axis. The orange crosses are bills that fail in the House, and the blue disks are bills that pass in the House. The error bar on each marker indicates the standard error. Our computation of closeness centrality incorporates an exponential decay with a half-life of 6 months. In the right panels, we plot the distributions of the relative differences between the closeness centralities of passed and failed bills. The horizontal axis gives the relative differences between the mean closeness centralities of bills that pass and fail in the House of Representatives. The vertical axis gives the counts of these relative differences.
Figure 6: Strength-centrality time series and the corresponding distributions of relative differences between the strength centralities of bills that pass and fail in the House of Representatives. In (a), we show the bill-level strength centrality that we calculate using the mean of the strength centralities of a bill’s cosponsors. In (b), we show the bill-level strength centrality that we calculate using the maximum of the strength centralities of a bill’s cosponsors. In the left panels, we show bill-level strength centrality for 10 years, which we aggregate into nonoverlapping 4-month periods. Each point is the mean strength centrality for bills in the associated 4-month period starting on the date on the horizontal axis. The orange crosses are bills that fail in the House, and the blue disks are bills that pass in the House. The error bar on each marker indicates the standard error. Our computation of strength centrality incorporates an exponential decay with a half-life of 6 months. In the right panels, we plot the distributions of the relative differences between the strength centralities of passed and failed bills. The horizontal axis gives the relative differences between the mean strength centralities of bills that pass and fail in the House of Representatives. The vertical axis gives the counts of these relative differences.
@@ A.5 Acknowledgements
We thank Martin Gilens and Zachary C. Steinert-Threlkeld for helpful comments.

### v2

@@ Appendix A Appendix
We now explore the influence of bill cosponsorship networks on bill-passage success using the standard centrality notions of closeness centrality (see Figure 6) and strength centrality (i.e., weighted degree centrality) (see Figure 7).
To compute closeness centrality and strength centrality [18], we use the largest connected component of a month-$t$ Congressional cosponsorship network and its corresponding adjacency matrix $D[t]_{\text{LCC}}$ (see Section 5.3). The closeness centrality of representative $i$ (i.e., node $i$) in month $t$ is
@@
where $N_{{\text{LCC}}}$ is the number of nodes in $D[t]_{\text{LCC}}$ and the entry $d_{ji}$ of $D[t]_{\text{LCC}}$ is the shortest-path distance from node $j$ to node $i$. The strength centrality of representative $i$ in month $t$ is
@@
Figure 6: Closeness-centrality time series and the corresponding distributions of relative differences between the closeness centralities of bills that pass and fail in the House of Representatives. In (a), we show the bill-level closeness centralities that we calculate using the mean of the closeness centralities of a bill’s cosponsors. In (b), we show the bill-level closeness centralities that we calculate using the maximum of the closeness centralities of a bill’s cosponsors. In the left panels, we show bill-level closeness centralities for 10 years, which we aggregate into nonoverlapping 4-month periods. Each point is the mean closeness centrality of bills in the associated 4-month period starting on the date on the horizontal axis. The values of the closeness centralities range between 0 and $1$. The orange crosses are bills that fail in the House, and the blue disks are bills that pass in the House. The error bar on each marker indicates the standard error. Our computation of closeness centralities incorporates an exponential decay with a half-life of 6 months. In the right panels, we plot the distributions of the relative differences between the closeness centralities of passed and failed bills. The horizontal axis gives the relative differences between the mean closeness centralities of bills that pass and fail in the House of Representatives. The vertical axis gives the counts of these relative differences.
Figure 7: Strength-centrality time series and the corresponding distributions of relative differences between the strength centralities of bills that pass and fail in the House of Representatives. In (a), we show the bill-level strength centralities that we calculate using the mean of the strength centralities of a bill’s cosponsors. In (b), we show the bill-level strength centralities that we calculate using the maximum of the strength centralities of a bill’s cosponsors. In the left panels, we show bill-level strength centralities for 10 years, which we aggregate into nonoverlapping 4-month periods. Each point is the mean strength centrality for bills in the associated 4-month period starting on the date on the horizontal axis. The values of the strength centralities range between 0 and $1$. The orange crosses are bills that fail in the House, and the blue disks are bills that pass in the House. The error bar on each marker indicates the standard error. Our computation of strength centralities incorporates an exponential decay with a half-life of 6 months. In the right panels, we plot the distributions of the relative differences between the strength centralities of passed and failed bills. The horizontal axis gives the relative differences between the mean strength centralities of bills that pass and fail in the House of Representatives. The vertical axis gives the counts of these relative differences.
@@ A.5 Acknowledgements
We thank Martin Gilens, Zachary C. Steinert-Threlkeld, Jennifer Nicoll Victor, and two anonymous referees for helpful comments.